# The Imperfective Paradox in Large Language Models

### v1

@@ 1 Introduction
Consider the sentence: “The carpenter was building a gazebo.”, as shown in Table 1. Logically, this description of an ongoing process does not entail that the gazebo was ever finished. The construction could have been abandoned or interrupted. This lack of entailment for goal-oriented events is known in formal semantics as the Imperfective Paradox (Dowty, 1977, 1979). In contrast, for activity verbs like run, the process itself constitutes the event Vendler (1957); Bennett and Partee (1978); Smith (1997): “The boy was running.” entails “The boy ran.”.
@@
    The Hegemony of Teleological Bias: Models exhibit a pervasive tendency to hallucinate completion for goal-oriented events. Crucially, this bias is robust enough to override explicit textual negation (e.g., predicting Entailment even when an action is interrupted), revealing a fundamental failure in contextual fidelity.
@@
    The Calibration Crisis in Prompting: Strategies like Chain-of-Thought or forced counterfactual non-endpoints can reduce hallucinated completions, but often over-correct, causing models to incorrectly reject valid entailments for atelic activities. Models struggle to adapt reasoning dynamically based on event type.
@@
    Dissociation between Representation and Reasoning: We find that the bias is not representational; embeddings successfully distinguish process from result. Instead, inference is dominated by prior expectations about goal completion. The failure occurs at decision time, not in representation.
Ultimately, our findings underscore that current LLMs struggle to resolve the Imperfective Paradox, operating as probabilistic predictors of narrative outcomes rather than faithful logical reasoners. This highlights that solving event semantics requires moving beyond surface-level alignment to structurally grounded aspectual reasoning.
@@ 2.1 Vendler’s Aspectual Classes: Telicity
  • Activities (Atelic): Events with no inherent endpoint (e.g., run, swim, push a cart). They are homogeneous; any part of the process counts as an instance of the same event.
  • Accomplishments (Telic): Events that proceed toward a specific terminus or goal (e.g., build a house, write a letter). Such events are only completed and receive canonical perfective interpretations, when the endpoint is reached.
@@ 2.3 Related Work: Interpreting and Modeling “Aspect” in NLP
Moving towards compositional semantics, Kober et al. (2020) introduced an NLI-based dataset to model aspect in context. With the rise of deep learning, Metheniti et al. (2022) analyzed transformer-based models, revealing that while they capture some aspectual signals, they exhibit significant biases towards surface features like tense and word order. Most recently, Ma (2024) evaluated large scale LLMs on lexical aspect identification and find LLMs are still short in understanding aspectual semantics. However, a critical gap remains in moving from aspectual identification to inference, specifically concerning whether LLMs respect the Imperfective Paradox or instead hallucinate event completion by failing to distinguish between goal-directed processes and their resulting states.
@@ 3 Dataset - ImperfectiveNLI
To evaluate the capability of LLMs in understanding the imperfective paradox and the teleological aspectual semantics, we introduce ImperfectiveNLI, a diagnostic NLI dataset in English.
@@
For each verb, we generated context-hypothesis pairs using Gemini-assisted Team et al. (2025) rewriting followed by strict human verification. The dataset is structured as a controlled minimal-pair experiment across four logical conditions, resulting in a total of $N=400$ examples.
@@
The carpenter was building a gazebo, but a storm destroyed it. The carpenter built a gazebo. Contradiction ( False)
@@ D Ambiguous Activity
Group A: Interrupted Accomplishment. This group serves as a counterfactual check. The premise describes a telic action in progress but includes an explicit cancellation clause (e.g., "…but a storm destroyed it"). This tests whether models can attend to explicit interruption of action and the negation of the result. The logical relation is Contradiction.
@@ 4.2 Prompts
Drawing on Dowty (1977)’s theory that the progressive aspect presupposes an “inertia world” (where events proceed normally), this prompt attempts to explicitly break this assumption. We arbitrarily instruct the model to first generate potential interruptions of events (with endpoint not achieving) before making a judgment.
@@ 4.3 Metrics
Table 3: Model performance under Zeroshot, CoT, DAP, and Counterfactual prompting strategies with Acc for A-D Groups, Teleological Bias Rate (TBR), and Aspectual Awareness Gap ($\Delta_{AA}$).
@@ 5.1 The Dominance of Teleological Bias for “If A Subject was V-ing, then the Subject V-ed” in Zero-shot
Contextual Neglect via Semantic Priming. Perhaps the most alarming finding is the resistance to explicit negation in Group A (Interrupted Accomplishment). Despite the textually grounded interruption (e.g., “…but a storm destroyed it”), most models exhibit a stark performance degradation. This stands in sharp contrast to Group B (Interrupted Activity), where models maintain high accuracy. This performance gap is semantically revealing: for activities, an interruption does not negate the event’s occurrence (the sub-interval property ensures that “stopped running” entails “ran”), whereas for accomplishments, interruption strictly precludes the result.
@@
The Outlier: Qwen2.5 shows exceptional performance. Qwen2.5 stands as a notable exception, maintaining a balanced performance with a TBR of $0.53$ and a healthy $\Delta_{AA}$ of $0.44$. Unlike its peers, Qwen demonstrates an emerging ability to suspend judgment in ambiguous contexts.
@@
Further evaluating three advanced prompting strategies illustrate a complex trade-off between mitigating teleological bias and maintaining semantic calibration across aspectual classes.
@@
DAP: The Knowledge-Application Gap. The DAP which introduces expert linguistic definition for the aspectual classes yields mixed results, exposing a critical gap between possessing linguistic knowledge and applying it. While supplying the linguistic rule improves Llama-3.1’s accuracy on Group C ($0.02\to 0.36$), it remains ineffective for Mistral ($0.06$). This persistence of bias implies that the teleological prior is not a misunderstanding of the task definition but a robust feature of the model’s probability distribution. As noted by Dziri et al. (2023), compositional reasoning in Transformers often degrades when it conflicts with high-frequency statistical patterns (i.e., the reporting bias) seen during pre-training.
@@
CoT: Reasoning-Induced Uncertainty. CoT prompting successfully reduces the TBR for most models, proving that allocating compute to intermediate reasoning helps suppress System-1 heuristics (Wei et al., 2022b). However, this comes at a cost. We observe a significant degradation in Group D (Activities), where accuracy drops substantially (e.g., Llama-3.1 $0.98\to 0.65$). Qualitative analysis reveals a phenomenon of “Over-reasoning” or unfaithful reasoning (Turpin et al., 2023). When forced to analyze “step-by-step”, models often hallucinate hypothetical interruptions for simple activities. [^3] See, e.g., the activity“running”, with model reasoning: “The premise describes a process in the present tense, indicating that the action of the athletes running on the track is ongoing or at least not explicitly completed. CoT appears to shift the model’s calibration towards excessive uncertainty, inadvertently disrupting the valid entailments of atelic verbs.
@@
Counterfactuals: The Calibration Crisis. The Counterfactual prompt represents the most aggressive intervention for arbitrary interruption. While highly effective at curing teleological bias (Group C Acc $\to 1.00$), it induces an obvious calibration bias. Accuracy on Group D collapses to near zero for Llama-3.1. We show the example of Llama-3.1 in Figure 1. The models swing from being “naively optimistic” (Zero-shot) to “paranoidly skeptical” (Counterfactual). The Counterfactual clearly over-instructed the models to follow the imperfective patterns of the telic verbs in Group C, resulting in missunderstanding for the atelic verbs in Group D. The models failed to dynamically adjust their reasoning strategy based on the (a)telicity of verbs, resulting in a wholesale rejection of progressive entailments. This shows one side the strong effect of prompts for the telic cases of the imperfective paradoxe, while it also highlights a lack of (aspectual) calibration (Kadavath et al., 2022; Kapoor et al., 2024): the ability to modulate uncertainty based on specific semantic features. Detailed analyses and visualization for other models are provided in Appendix E.
@@
Figure 1: The trade-off in prompting (Llama-3.1). As prompts become more aggressive in targeting bias (Zero-shot $\to$ Counterfactual), accuracy on Group C improves significantly (Pearson $r=1.00,p=0.00$), and accuracy on Group D collapses also, only unsteadily (r=0.91,p=0.09r=-0.91,p=0.09). This demonstrates the calibration bias.
@@
Summary on Different Prompts. In summary, current LLMs struggle to balance the Imperfective Paradox. They oscillate between two extremes: naive teleology (Zero-shot) where everything is completed, and paranoid skepticism (Counterfactual) where nothing is ever certain (see the trends in Figure 1). This inability to decouple telicity from entailment highlights a critical gap in the true aspectual reasoning capabilities of LLMs.
@@ 5.3 Impact of Model Scale
The 1.5B model exhibits a complete inability to handle the Imperfective Paradox. With a TBR of $1.00$ and a $\Delta_{AA}$ of $0.00$, it treats every progressive verb as an entailed completion. With increasing model scales, especially the 32B model demonstrates a dramatic “phase shift” (Wei et al., 2022a): Accuracy on Group C surges to $0.91$, and the $\Delta_{AA}$ peaks at $0.83$. This suggests that robust aspectual distinction, the ability to selectively inhibit the entailment for telic verbs while maintaining it for atelic ones, is an emergent ability. Although the 72B model maintains strong performance, $\Delta_{AA}$ decreases slightly to 0.81, suggesting a potential saturation or optimization bottleneck rather than further qualitative gains. Overall, these results indicate that aspectual understanding improves with scale and exhibits non-linear emergence characteristics rather than linear scaling.
@@ 5.4 Fine-grained Analysis on Verb Semantics
Do all telic events trigger the same degree of teleological bias? To answer this, we disaggregate the performance on Group A and Group C based on the four semantic sub-classes defined in §3: Creation, Consumption, Change of State, and Motion to Goal. Figure 3 visualizes the average accuracy across all evaluated models in the zero-shot setting. Detailed results for each model as well as for other prompt settings can be found in Appendix E.

### v2

@@ 1 Introduction
Consider the following sentence: “The carpenter was building a gazebo”, as shown in Table 1. Logically, this description of an ongoing process does not entail that the gazebo was ever finished. The construction could have been abandoned or interrupted. This lack of entailment for telic (goal-oriented) events is known in formal semantics as the Imperfective Paradox (Dowty, 1977, 1979). In contrast, for atelic verbs like “run”, the process itself constitutes the event Vendler (1957); Bennett and Partee (1978); Smith (1997): “The boy was running” entails “The boy ran”.
@@
    Teleological Bias Dominates Model Inference: Models exhibit a pervasive tendency to hallucinate completion for goal-oriented events. Crucially, this bias is robust enough to override explicit textual negation (e.g., predicting completion even when an action is interrupted), revealing a fundamental failure in contextual fidelity.
@@
    Prompting Interventions Trigger a Calibration Crisis: Strategies like Chain-of-Thought or forced counterfactual non-endpoints can reduce hallucinated completions, but often over-correct, causing models to incorrectly reject valid entailments for atelic activities. Models struggle to adapt reasoning dynamically based on event type.
@@
    Representation and Reasoning Are Dissociated: We find that the bias is not representational; embeddings successfully distinguish process (progressive form) from result (simple past form). Instead, inference is dominated by prior expectations about goal completion. The failure is likely driven by predictive priors during decoding, rather than an inherent inability of Transformer architectures to encode the aspectual distinction.
Ultimately, our findings underscore that current LLMs struggle to resolve the Imperfective Paradox, operating as probabilistic predictors of narrative outcomes rather than faithful logical reasoners. Bridging this gap requires not just better prompting, but a deeper alignment between model behavior and the formal logic of event semantics.
@@ 2.1 Vendler’s Aspectual Classes: Telicity
  • Activities (Atelic): Events with no inherent endpoint (e.g., “run”, “swim”, “push a cart”). They are homogeneous; any part of the process counts as an instance of the same event.
  • Accomplishments (Telic): Events that proceed toward a specific terminus or goal (e.g., “build a house”, “write a letter”). Such events are only completed and receive canonical perfective interpretations, when the endpoint is reached.
@@ 2.3 Related Work: Interpreting and Modeling “Aspect” in NLP
Moving toward compositional semantics, Kober et al. (2020) introduced an NLI-based dataset to model aspect in context. Later work showed that transformer models capture some aspectual cues but rely heavily on surface features such as tense and word order (Metheniti et al., 2022). More recently, Ma (2024) found that large-scale LLMs still struggle with lexical aspect identification. Closely related are two studies on the Imperfective Paradox Pruś et al. (2024, 2025), which focus primarily on modeling human judgments of progressive versus simple past forms, treating LLM behavior as a secondary comparison. As such, neither study provides a systematic diagnostic evaluation of LLM aspectual inference. Our work directly addresses this gap, offering the first controlled investigation of whether LLMs respect the Imperfective Paradox or instead hallucinate event completion by conflating goal-directed processes with their result states.
@@ 3 Dataset - ImperfectiveNLI
To evaluate LLMs’ ability to reason about the imperfective paradox and teleological aspectual semantics, we introduce ImperfectiveNLI, a diagnostic NLI dataset in English.
@@
For each verb, we generated context-hypothesis pairs using Gemini-assisted rewriting Team et al. (2025), followed by strict human verification. The dataset is structured as a controlled minimal-pair experiment across four logical conditions, resulting in a total of $N=400$ examples.
@@
The carpenter was building a gazebo, but a storm destroyed the frame before the roof was on. The carpenter built a gazebo. Contradiction ( False)
@@ D Ambiguous Activity
Imperfective Logical Conditions.
Group A: Interrupted Accomplishment. This group serves as a counterfactual check. The premise describes a telic action in progress but includes an explicit cancellation clause (e.g., “…but a storm destroyed the frame before the roof was on.”). This tests whether models can attend to explicit interruption of action and the negation of the result. The logical relation is Contradiction.
@@ 4.2 Prompts
Drawing on Dowty (1977)’s theory that the progressive aspect presupposes an “inertia world” (where events proceed normally), this prompt attempts to explicitly break this assumption. We explicitly instruct the model to first generate three plausible real-world scenarios in which the event is interrupted and its endpoint is not reached, before making a judgment.
@@ 4.3 Metrics
Note that $\Delta_{AA}$ is most interpretable when $\text{ACC}_{D}$ is high; a low $\Delta_{AA}$ driven by a low $\text{ACC}_{D}$ rather than a high $\text{TBR}_{C}$ reflects a distinct failure mode, namely, an inability to recognize valid entailments for atelic verbs, rather than teleological overgeneralization.
@@
Table 3: Model performance under Zero-shot, CoT, DAP, and Counterfactual prompting strategies with Acc for A-D Groups, Teleological Bias Rate (TBR), and Aspectual Awareness Gap ($\Delta_{AA}$).
@@ 5.1 The Dominance of Teleological Bias for “If A Subject was V-ing, then the Subject V-ed” in Zero-shot
At first glance, models appear competent in aspectual reasoning, achieving near-perfect accuracy on Group D for ambiguous Activity (atelic verbs) (e.g., Llama-3.1: $0.98$, Mistral: $1.00$). However, this high performance is deceptive. When contrasted with the catastrophic failure on Group C for ambiguous Accomplishment (telic verbs) (e.g., Llama-3.1: $0.02$), a different picture emerges. The near-zero $\Delta_{AA}$ for most models indicates that they do not genuinely understand the Sub-interval Property of the atelic verbs (Bennett and Partee, 1978). Instead, they likely employ a shallow heuristic: “If Subject was V-ing, then Subject V-ed”, regardless of verbs in the Zero-shot setting. This aligns with findings by the NLI work McCoy et al. (2019), suggesting that the models might often rely on lexical overlap strategies rather than deep semantic compositionality.
Contextual Neglect via Semantic Priming.
Perhaps the most alarming finding is the resistance to explicit negation in Group A (Interrupted Accomplishment). Despite the textually grounded interruption (e.g., “…but a storm destroyed it”), most models exhibit a stark performance degradation. This stands in sharp contrast to Group B (Interrupted Activity), where models maintain high accuracy. This performance gap is semantically revealing: for activities, an interruption does not negate the event’s occurrence (the sub-interval property ensures that “stopped running” entails “ran”), whereas for accomplishments, interruption strictly precludes the result.
@@
The Outlier: Qwen2.5 shows exceptional performance.
Qwen2.5 stands as a notable exception, maintaining a balanced performance with a TBR of $0.53$ and a healthy $\Delta_{AA}$ of $0.44$. Unlike its peers, Qwen demonstrates an emerging ability to suspend judgment in ambiguous contexts.
@@
Evaluating three advanced prompting strategies reveals a consistent trade-off between mitigating teleological bias and maintaining semantic calibration across aspectual classes.
DAP: The Knowledge-Application Gap.
@@
The DAP which introduces expert linguistic definition for the aspectual classes yields mixed results, exposing a critical gap between possessing linguistic knowledge and applying it. While supplying the linguistic rule improves Llama-3.1’s accuracy on Group C ($0.02\to 0.36$), it remains ineffective for Mistral ($0.06$). This persistence of bias implies that the teleological prior is not a misunderstanding of the task definition but a robust feature of the model’s probability distribution. As noted by Dziri et al. (2023), compositional reasoning in Transformers often degrades when it conflicts with high-frequency statistical patterns (i.e., the reporting bias) seen during pre-training.
CoT: Reasoning-Induced Uncertainty.
@@
CoT prompting successfully reduces the TBR for most models, proving that allocating compute to intermediate reasoning helps suppress System-1 heuristics (Wei et al., 2022b). However, this comes at a cost. We observe a significant degradation in Group D (Activities), where accuracy drops substantially (e.g., Llama-3.1 $0.98\to 0.65$). Qualitative analysis reveals a phenomenon of “over-reasoning” or unfaithful reasoning (Turpin et al., 2023). When forced to analyze “step-by-step”, models often hallucinate hypothetical interruptions for simple activities. [^3] See, e.g., the activity“running”, with model reasoning: “The premise describes a process in the present tense, indicating that the action of the athletes running on the track is ongoing or at least not explicitly completed. CoT appears to over-generalize the progressive-as-incomplete heuristic: when forced to reason step-by-step, models often impose hypothetical interruptions even on atelic verbs, inadvertently suppressing valid entailments.
Counterfactuals: The Calibration Crisis.
@@
The Counterfactual prompt represents the most aggressive intervention for arbitrary interruption. While highly effective at curing teleological bias (Group C Acc $\to 1.00$), it induces an obvious calibration bias. Accuracy on Group D collapses to near zero for Llama-3.1. We show the example of Llama-3.1 in Figure 1. The models swing from being “naively optimistic” (Zero-shot) to “paranoidly skeptical” (Counterfactual). The Counterfactual clearly over-instructed the models to follow the imperfective patterns of the telic verbs in Group C, resulting in misunderstanding for the atelic verbs in Group D. The models failed to dynamically adjust their reasoning strategy based on the (a)telicity of verbs, resulting in a wholesale rejection of progressive entailments. This highlights a lack of (aspectual) calibration (Kadavath et al., 2022; Kapoor et al., 2024): while the counterfactual prompt successfully suppresses teleological bias for telic verbs, models fail to modulate this skepticism based on verb class, applying it indiscriminately to atelic verbs as well. Detailed analyses and visualization for other models are provided in Appendix E.
@@
Figure 1: The trade-off in prompting (Llama-3.1). As prompts become more aggressive in targeting bias (Zero-shot $\to$ Counterfactual), accuracy on Group C improves significantly (Pearson $r=1.00,p=0.00$), while accuracy on Group D declines, though less consistently (r=0.91,p=0.09r=-0.91,p=0.09). This trade-off illustrates the calibration bias.
Summary on Different Prompts.
@@
In summary, current LLMs struggle to balance the Imperfective Paradox. They oscillate between two extremes: naive teleology (Zero-shot) where everything is completed, and paranoid skepticism (Counterfactual) where nothing is ever certain (see the trends in Figure 1). This inability to decouple telicity from entailment highlights a critical gap in the true aspectual reasoning capabilities of LLMs.
@@ 5.3 Impact of Model Scale
The 1.5B model exhibits a complete inability to handle the Imperfective Paradox. With a TBR of $1.00$ and a $\Delta_{AA}$ of $0.00$, it treats every progressive verb as an entailed completion. As model scale increases, a dramatic “phase shift” (Wei et al., 2022a) is observed, especially at the 32B level: Accuracy on Group C surges to $0.91$, and the $\Delta_{AA}$ peaks at $0.83$. This suggests that robust aspectual distinction, the ability to selectively inhibit the entailment for telic verbs while maintaining it for atelic ones, is an emergent ability. Although the 72B model maintains strong performance, $\Delta_{AA}$ decreases slightly to 0.81, suggesting a potential saturation or optimization bottleneck rather than further qualitative gains. Overall, these results indicate that aspectual understanding improves with scale and exhibits non-linear emergence characteristics rather than linear scaling.
@@ 5.4 Fine-grained Analysis on Verb Semantics
Do all telic verb classes trigger the same degree of teleological bias? To answer this, we disaggregate the performance on Group A and Group C based on the four semantic sub-classes defined in §3: Creation, Consumption, Change of State, and Motion to Goal. Figure 3 visualizes the average accuracy across all evaluated models in the zero-shot setting. Detailed results for each model as well as for other prompt settings can be found in Appendix E.